Fred has returned for his next visit.

Client: Fred

Age: 56

Height: 5 feet 10 inches                            Weight: 270 pounds         

BMR calculations:

Males BMR= 66+(6.23 x weight in pounds) +(12.7 x height in inches) -(6.8 x age)

Physical Activity (PA) Estimates

1.2 = little to no exercise

1.375 = light exercise 1-3 days per week

1.55 = moderate exercise 3-5 days per week

1.725 = hard exercise 6-7 days per week

1.9 = hard daily exercise/job

1. Today you want to determine the amount of energy (calories) that Fred should be ingesting daily to maintain his current weight. Using the BMR formula, determine Fred’s BMR.

2. To determine individual energy requirements, multiply the BMR x the appropriate activity level. This provides the number of calories required daily to maintain the current weight.
3. Using a BMI chart, Fred would be considered? ___________________

4. The medical records chart indicates that Fred’s PCP recommends that Fred lose weight. Review the three components of a weight loss program with Fred.
Answer #1

Client:-.Fred

Age:-56

Height:-5 feet 10 inches (70 inches)

weight:- 270 pounds

1)Formula of Male BMR =66+ (6.23 × weight in pounds) + (12.7× height in inches) - (6.8× Age)

Fred BMR= 66 + (6.23 × 270) + (12.7 × 70) - (6.8× 56)

= 66 + 1682.1 + 889 - 380.8

= 2637.1 - 380.8

= 2256.3 kcal

2)Energy Requirement= BMR× the appropriate activity level

Suppose Fred do moderate activity daily ,then energy requirement= 2256.3 × 1.55= 3497.265 Kcal

3)BMI= Weight(Pound)×703/Height(inches) square

Fred BMI= 270× 703/70× 70

= 189810/4900= 38.73

Normal BMI- 18.5 - 25, BMI - 25- 30 (overweight), BMI>30( obese)

Fred would be considered obese( BMI- 38.73)

4)Three components of weight loss program with Fred are as follows:-

a)Good Nutrition is the main weight loss program.Good nutrition means choosing whole foods and avoiding processed foods as they are often high in sugar,fat,salt and calories but low in nutrients.Whole foods such as fresh leafy vegetables,fruits,eggs,lean meats are nutrient dense.To loss weight, Fred would need to consume fewer calories than he expends so it is important to be aware of calories and portion size.

b)Physical activity:--Fred would choose an activity that he enjoys and can do regularly.He should do 30 to 40 minutes daily either walking, swimming or rowing as physical activity burns extra unwanted fats.

C)Modification in behavior and emotions:-

Alongwith dietary and physical activity , behavior and emotions related to weight management are also essential to successful weight loss.Fred should boost his self esteem and support his weight loss efforts.